40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 2591     Entry time: Mon Jun 21 10:59:41 2021
Author: Radhika 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: Chamber updates 6/16-6/18 

[Radhika, Stephen]

Background: There are currently 4 available RTDs for the large cryo chamber (henceforth named Megastat). We originally had 1 for the workpiece, 1 for the heater, 1 for the inner shield, and 1 for the outer shield. During the last sessions with the chamber open, the RTDs were moved around. Now there is 1 for the workpiece, 1 for the baseplate (bottom lid), one for the outer shield, and 1 for the cold head. These locations are marked in the diagram (attachment 9). 

On Wednesday 6/16, we opened up the chamber and removed the shields and coldplate. We added foil to the inner surface of the bottom lid (attachment 10) and attached an RTD there (we forgot to take a picture). Stephen made a new foil collar, and we decided to push it against the chamber walls as a better alternative to having the foil touch the coldplate (attachment 6). We added an extra layer of aluminized mylar to wrap the copper braid, and we fed another RTD through the braid tube to be attached to the coldhead. We lastly placed the coldplate and shields back in place.

On Friday 6/18, I attached the remaining RTDs (coldhead, outer shield, workpiece). We decided to remove the aluminum foil previously covering the coldhead, due to of fear of shorting to the tube wall. I taped several pieces of mylar together to cover the coldhead and insulate it from the wall (attachments 1-3). I placed the mylar contraption into the T (attachment 4) and then closed the bottom flange. I placed back the shield lids and the main lid of the chamber (attachments 7-8). I tried to pump down, but the pressure was stabilizing on the order of 1e-1 torr. Unsure of why the pressure wasn't decreasing, I turned off the pump and left for the day.

Today 6/21, I touched base with Stephen and we realized I forgot to replace the copper gasket in the bottom flange of the T. I then unscrewed and replaced the gasket. I pumped down and the pressure reached on the order of 1e-4 torr, so I proceeded with cooling. After a few hours I could see that the cold head RTD was reading a temperature around 180K. We will extract and plot cooling data late tomorrow or Wednesday.

Lastly, a chamber diagram is attached (attachment 9). The 4 RTD locations are marked by a red 'x'. The chamber components are numbered in blue (detailed below):

1. Workpiece

2. Heater

3. Inner shield

4. Outer shield

5. Copper braid (wrapped with mylar)

6. Cold head

7. Coldplate

8. Baseplate (bottom lid)

Attachment 1: IMG_1735.jpeg  1.839 MB  Uploaded Mon Jun 21 12:14:04 2021  | Hide | Hide all
IMG_1735.jpeg
Attachment 2: IMG_1738.jpeg  2.076 MB  Uploaded Mon Jun 21 12:14:55 2021  | Hide | Hide all
IMG_1738.jpeg
Attachment 3: IMG_1737.jpeg  2.020 MB  Uploaded Mon Jun 21 12:15:09 2021  | Hide | Hide all
IMG_1737.jpeg
Attachment 4: Photo_on_6-18-21_at_15.36.jpeg  69 kB  Uploaded Mon Jun 21 12:15:23 2021  | Hide | Hide all
Photo_on_6-18-21_at_15.36.jpeg
Attachment 5: Photo_on_6-18-21_at_16.32.jpeg  92 kB  Uploaded Mon Jun 21 12:15:34 2021  | Hide | Hide all
Photo_on_6-18-21_at_16.32.jpeg
Attachment 6: IMG_8947.jpeg  157 kB  Uploaded Mon Jun 21 12:17:57 2021  | Hide | Hide all
IMG_8947.jpeg
Attachment 7: Photo_on_6-18-21_at_16.33.jpeg  99 kB  Uploaded Mon Jun 21 12:18:07 2021  | Hide | Hide all
Photo_on_6-18-21_at_16.33.jpeg
Attachment 8: Photo_on_6-18-21_at_16.33_#2.jpeg  109 kB  Uploaded Mon Jun 21 12:18:15 2021  | Hide | Hide all
Photo_on_6-18-21_at_16.33_#2.jpeg
Attachment 9: New_Note.png  363 kB  Uploaded Mon Jun 21 12:19:55 2021  | Hide | Hide all
New_Note.png
Attachment 10: IMG_8973.jpeg  437 kB  Uploaded Wed Jun 23 16:50:55 2021  | Hide | Hide all
IMG_8973.jpeg
ELOG V3.1.3-